OpenStudy (anonymous):

The formula for calculating the amount of money returned for an initial deposit into a bank account or CD (certificate of deposit) is given by A=p(1+r/n)^nt A is the amount of the return. P is the principal amount initially deposited. r is the annual interest rate (expressed as a decimal). n is the number of compound periods in one year. t is the number of years. Carry all calculations to six decimals on each intermediate step, then round the final answer to the nearest cent. Suppose you deposit $3,000 for 6 years at a rate of 7%. a) Calculate the return (A) if the bank compounds semi-annually. B. Calculate the return (A) if the bank compounds monthly. C. If a bank compounds continuously, then the formula used is A= Pe^rt where e is a constant and equals approximately 2.7183. Calculate A with continuous compounding. Round your answer to the nearest cent. Answer: Show work in this space

OpenStudy (anonymous):

a) Calculate the return (A) if the bank compounds semi-annually. @leojones77 Can you plug the numbers into A=p(1+r/n)^nt with n = 2

OpenStudy (anonymous):

Similarly with n = 12 for B) compounds monthly.

OpenStudy (anonymous):

Can you show me how I would set this problem up so that I can work it out please

OpenStudy (anonymous):

You already have the formula and all the data ready, just plug numbers in !

OpenStudy (anonymous):

A= P (1 + r/n ) ^nt with n = 2 = 3,000 ( 1 + .07/2 )^(2*6)

OpenStudy (anonymous):

thanks